Hi I was wondering. I need help with my help with my math hw. I'm have to solve a problem. Solve each equation using substitution. Y =4/3x-2. Y=2/3x
Answer by Alan3354(69443)

Solve each equation using substitution. Y =4/3x-2. Y=2/3x
----------
Sub for y
(4/3)x - 2 = (2/3)x
Multiply by 3
4x - 6 = 2x
2x = 6
x = 3
----
y = 2